# DSMoE: Matrix-Partitioned Experts with Dynamic Routing for Computation-Efficient Dense LLMs

### v2

@@ Abstract
Minxuan Lv 1,2111These authors contributed equally to this work., Zhenpeng Su1,2,3111These authors contributed equally to this work., Leiyu Pan4,Yizhe Xiong5,Zijia Lin3,5222Corresponding authors.,Hui Chen5222Corresponding authors.,Wei Zhou1,2 Jungong Han5,Guiguang Ding5,Cheng Luo3,Di Zhang3,Kun Gai3,Songlin Hu1,$2$Corresponding authors. 1Institute of Information Engineering, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Kuaishou Technology, 4Tianjin University, 5Tsinghua University {lvminxuan,husonglin}@iie.ac.cnhuichen@tsinghua.edu.cn {suzhenpeng,linzijia}@kuaishou.com
@@ 1 Introduction
At the algorithmic level, approaches to model efficiency optimization generally follow two paradigms: post-training compression and acceleration of dense models, or training of Mixture of Experts (MoE) architectures. While compression methods like pruning achieve efficiency through permanent parameter removalAshkboos et al. (2024); Ma et al. (2023); Frantar and Alistarh (2023), they may discard valuable knowledge and lack flexibility in handling inputs of varying complexity. Conversely, although effective, MoE approaches—whether trained from scratchFedus et al. (2022); Dai et al. (2024); Liu et al. (2024) or warm-started from dense modelsJiang et al. (2024)—cannot reduce the computational cost of the original dense model. Given that the most widely used and effective foundation models still maintain dense architectures (such as LLaMATouvron et al. (2023), QwenBai et al. (2023)), there is a crucial need to optimize these widely-adopted dense models for better efficiency without sacrificing their performance
To address these limitations, we propose DSMoE, which achieves model sparsification by partitioning pre-trained FFN layers into smaller computational blocks. Unlike existing approaches, DSMoE preserves all model parameters, eliminating the risk of knowledge loss, while achieving lower computational overhead than the original dense model through dynamic routing mechanisms. This design maintains the model’s complete knowledge while improving computational efficiency.
@@ 2 Related Work
In recent years, there has been growing interest in exploring sparse computation in large language models. Mixture-of-Experts (MoE) represents a pioneering approach that demonstrates how sparse activation can effectively balance model capacity and computational efficiency. In MoE architectures, only a subset of FFN modules (experts) are activated for each input token Fedus et al. (2022); Lepikhin et al. (2021); Huang et al. (2024). This foundational idea of conditional computation has inspired various innovations in expert activation strategies. Some works explore heterogeneous expert architectures Sun et al. (2024) or introduce zero-computation experts Jin et al. (2024) to further optimize computational efficiency. These advances in MoE architectures demonstrate the potential of sparse computation and motivate our exploration of applying similar principles within individual FFN layers.
@@ 4 Method
Although our method is termed DSMoE, its training approach differs from traditional MoE methods such as Switch Transformer Fedus et al. (2022) and DeepSeeKMoE Dai et al. (2024). Our objective is to achieve sparsity through partitioning pre-trained models, where each expert inherits a distinct portion of the original model’s knowledge. Our approach is based on the principle that the model should learn to selectively utilize different aspects of pre-trained knowledge based on input complexity, rather than routing tokens among independently trained experts. To implement this insight, we present our method in three modules.
@@ 4.1 FFN Partitioning
The widespread adoption of MoE architectures inspires our exploration of sparsity in FFN layers, suggesting that different parts of computation can be dynamically activated based on input patterns. Previous work has further revealed that FFN layers essentially operate as key-value memories, where different portions of the layer specialize in detecting and processing distinct input patternsGeva et al. (2020). Building on these insights, we propose to directly partition pre-trained FFN layers. As shown in Equation 5, we partition the matrices $\mathbf{U}$, $\mathbf{V}$, and $\mathbf{W}$ into $n$ groups along the intermediate dimension, where each group can be viewed as an “expert" that inherits a portion of the original transformation capabilities. When summing all expert outputs, this partitioned form yields identical results to the original FFN computation:
@@
We can structurally split the original FFN layer matrix into multiple small FFN matrices. To enable dynamic expert activation based on input, we employ a gating network that determines which experts should be activated. The expert’s output is propagated to the subsequent layer only when the corresponding gating activation value exceeds a certain threshold $\tau$. This can be formulated as:
@@
Model Configuration Params Activated Params FLOPs PPL ($\downarrow$)
LLaMA-1B d=2048, D=8192 1.24B 1.24B 2.53 T 5.67
LLaMA-7B d=4096, D=11008 6.74B 6.74B 13.53 T 3.40
@@
LLM-Pruner-channel d=1215, D=8192 889M 889M 1.50 T 7.51
LLM-Pruner-block d=2048, D=3896.4 735M 735M 1.50 T 7.46
SparseGPT d=2048, D=8192 1.24B 1.24B 2.53 T 9.82
MoE d=2048, D=1024    ×$\times 8$, topK=3 1.24B 736M 1.50 T 7.45
DSMoE(ours) d=2048, D=1024    ×$\times 8$ 1.24B 735M 1.50 T 7.41
@@
LLM-Pruner-channel d=2401,   D=11008 3.95B 3.95B 7.93 T 4.01
LLM-Pruner-block d=11008, D=6256.5 3.94B 3.94B 7.93 T 4.01
SparseGPT d=4096,    D=11008 6.74B 6.74B 13.53 T 3.96
MoE d=2048,    D=1376    ×$\times 8$, topK=3 6.74B 3.98B 7.99 T 4.12
DSMoE(ours) d=2048,   D=1376    ×$\times 8$ 6.74B 3.93B 7.91 T 3.91
Table 1: Results of perplexity (PPL) across different language models. The bold values indicate the best-performing method among various acceleration approaches. The Configuration column describes the specific model architecture, where $d$ represents the hidden dimension, D denotes the expansion dimension in FFN layers (for LLM-Pruner-block method, this represents the average value), $\times$ n indicates the use of n parallel FFN layers, and topK specifies the number of activated experts per layer in the MoE architecture. The Params column shows the total number of model parameters, while Activated Params indicates the average number of parameters activated during inference. FLOPs (Floating Point Operations) represents the average number of floating-point operations per sample.
@@ 4.2 Straight-Through Estimator
This gradient blocking prevents non-activated experts from receiving training signals, leading to a “dead expert" problem where these experts become permanently inactive. Unlike traditional MoE models that train experts from scratch, our experts inherit pre-trained knowledge that we wish to preserve and adapt. To address this issue, we employ the straight-through estimator technique, which allows gradient flow through non-activated experts while maintaining thresholded activation during the forward pass:
$S(x)=sg(G(x))+x-sg(x)$ (10)
@@
where the operator “$sg(\cdot)$" is the “stop gradient" operator to prevent gradient back propagation. The partial derivatives for experts and their gates below the threshold are as follows. Let:
$a_{i}=\text{act}(\hat{h}_{t}^{l}\mathbf{W}_{i})$ (12)
$a^{\prime}_{i}=\text{act}^{\prime}(\hat{h}_{t}^{l}\mathbf{W}_{i})$
$g_{i}=\sigma(\hat{h}_{t}^{l}\mathbf{Y}_{i})$
$u_{i}=\hat{h}_{t}^{l}\mathbf{U}_{i}$
@@
$\frac{\partial h_{t}^{l}}{\partial\mathbf{Y}_{i}}=(\hat{h}_{t}^{l})^{\top}\cdot(o_{i}\cdot\sigma^{\prime}(\hat{h}_{t}^{l}\mathbf{Y}_{i}))$ (16)
The gradient dynamics reveal an important property: since $\sigma^{\prime}(\hat{h}_{t}^{l}\mathbf{Y}_{i})>0$, an expert that produces meaningful output $o_{i}$ for an input $(\hat{h}_{t}^{l})^{\top}$ will receive gradients that increase its activation probability for similar inputs in future iterations, regardless of its current activation status. This adaptive mechanism ensures that experts can learn to specialize in processing specific input patterns while maintaining their inherited knowledge from pre-training.
@@ 4.3 Sparse Loss
$\mathcal{L}=\mathcal{L}_{\text{LM}}+\lambda\mathcal{L}_{\text{sparse}}$ (17)
where $\mathcal{L}_{\text{sparse}}$ denotes the sparsity loss term, which we abbreviate as $\mathcal{L}{\text{s}}$ in subsequent equations. The hyperparameter $\lambda$ controls the strength of sparsity regularization, with larger values encouraging sparser activation patterns.
$\mathcal{L}=\mathcal{L}_{\text{LM}}+\frac{\lambda}{LN}\sum_{l=1}^{L}\sum_{n=1}^{N}\mathcal{L}_{s}(G(\sigma(\hat{h}^{l}_{t}\mathbf{Y}_{n})))$ (18)
@@
$\mathcal{L}=\mathcal{L}_{LM}+\frac{\lambda}{LN}\sum_{l=1}^{L}\sum_{n=1}^{N}G(\sigma(\hat{h}^{l}_{t}\mathbf{Y}_{n}))$ (19)
@@ 5 Experiments
Model Hellaswag LAMBADA PIQA SIQA StoryCloze Wino GSM8K TriviaQA WebQs NatrualQs
@@ 5.2 Experimental Setup
We evaluate DSMoE on two pre-trained models of different scales: Llama-7B [^1] https://huggingface.co/meta-llama/Llama-2-7b and Llama-1B [^2] https://huggingface.co/meta-llama/Llama-3.2-1B. For our method’s hyperparameters, we simply set the activation threshold $\tau=0.5$ and the sparsity regularization coefficient $\lambda=1.0$.
We compare our approach with several baselines: the channel-wise and block-wise methods from LLM-Pruner (a structured pruning approach), and SparseGPT (an unstructured pruning method). To ensure fair comparison, we first measure the FLOPs of our trained model, then estimate the pruning ratio for baseline methods to maintain a slightly higher FLOPs than our method. The FLOPs metric directly corresponds to the number of parameters involved in computation, providing a standardized measure of computational efficiency.
Additionally, we explore an alternative approach by applying the same FFN partitioning scheme but training it as a traditional MoE architecture (with fixed expert selection and standard MoE training objectives) to investigate whether the conventional MoE framework better accommodates the warm-starting paradigm.
@@ 5.3 Main Results
We first present the model’s perplexity on the validation set. Following previous workTouvron et al. (2023); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we then evaluate the model’s performance on downstream benchmarks, which includes zero-shot accuracy testing on HellaSwagZellers et al. (2019), LAMBADAPaperno et al. (2016), SIQASap et al. (2019), PIQABisk et al. (2020), StoryClozeMostafazadeh et al. (2016), and WinograndeSakaguchi et al. (2021). Additionally, we conduct 5-shot evaluation measuring exact match performance on TriviaQAJoshi et al. (2017), WebQuestions (WebQs)Berant et al. (2013), GSM8KCobbe et al. (2021), and Natural Questions (NaturalQs)Kwiatkowski et al. (2019).
@@ 5.3.1 Perplexity Results
Table 1 presents the perplexity results of the baseline dense model and its pruned, sparsified variants. The results demonstrate that DSMoE consistently outperforms baseline models under equivalent activation constraints. Since SparseGPT acceleration requires specific pruning ratios and hardware support, we conducted our comparative analysis only on models with equivalent parameter pruning levels. Our experimental results indicate that DSMoE achieves superior efficiency compared to static parameter pruning. Furthermore, DSMoE exhibits better performance than fixed-activation methods like MoE, which can be attributed to the fact that knowledge from all experts contributes to the model’s learning process, enabling it to develop the ability to flexibly select activations based on input. Additionally, DSMoE exhibits distinctive feature processing capabilities, learning layer-specific activation patterns that naturally emerge from the input complexity. We will examine these emergent patterns in detail in the analysis section.

### v3

@@ Abstract
Minxuan Lv 1,2111These authors contributed equally to this work., Zhenpeng Su1,2,3111These authors contributed equally to this work., Leiyu Pan4,Yizhe Xiong5,Zijia Lin3,5222Corresponding authors.
This work was supported by National Natural Science Foundation of China (Nos. 62271281, 62441235, 62525103). It was also sponsored by CCF-Kuaishou Large Model Explorer Fund (No. CCF-Kuaishou 2024001).
,Hui Chen5222Corresponding authors.
This work was supported by National Natural Science Foundation of China (Nos. 62271281, 62441235, 62525103). It was also sponsored by CCF-Kuaishou Large Model Explorer Fund (No. CCF-Kuaishou 2024001).
,Wei Zhou1,2
Jungong Han5,Guiguang Ding5,Cheng Luo3,Di Zhang3,Kun Gai3,Songlin Hu1,2222Corresponding authors.
This work was supported by National Natural Science Foundation of China (Nos. 62271281, 62441235, 62525103). It was also sponsored by CCF-Kuaishou Large Model Explorer Fund (No. CCF-Kuaishou 2024001).
1Institute of Information Engineering, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Kuaishou Technology, 4Tianjin University, 5Tsinghua University {lvminxuan,husonglin}@iie.ac.cnhuichen@tsinghua.edu.cn {suzhenpeng,linzijia}@kuaishou.com
@@ 1 Introduction
At the algorithmic level, approaches to model efficiency optimization generally follow two paradigms: post-training compression and acceleration of dense models, or training of Mixture of Experts (MoE) architectures. While compression methods like pruning achieve efficiency through permanent parameter removalAshkboos et al. (2024); Ma et al. (2023); Frantar and Alistarh (2023), they may discard valuable knowledge and lack flexibility in handling inputs of varying complexity. Conversely, though MoE approaches effectively expand model capacityFedus et al. (2022); Dai et al. (2024); Liu et al. (2024), traditional MoE typically employs fixed activation patterns where each token can only access a predetermined number of experts, lacking the ability to dynamically adjust computation based on input complexity. Given that the most widely used and effective foundation models still maintain dense architectures (such as LLaMATouvron et al. (2023), QwenBai et al. (2023)), we face a critical challenge: how to achieve truly input-adaptive computation while preserving pre-trained knowledge, allowing models to dynamically adjust activated parameters according to varying input complexity, thereby reaching an optimal balance between computational efficiency and model performance.
To address this challenge, we propose DSMoE, a novel approach that partitions pre-trained FFN layers into computational blocks and introduces dynamic routing mechanisms. DSMoE fundamentally differs from existing methods by preserving the original model parameters and reorganizing them into expert networks, while incorporating adaptive routing mechanisms that enable dynamic expert activation based on input complexity, rather than fixed activation strategies. Through straight-through estimator and sparsity loss design, DSMoE enables the model to autonomously learn sparse expert activation patterns, achieving computational resource allocation for inputs of varying complexity.
@@ 2 Related Work
The Mixture of Experts architecture is recognized as a promising approach for model sparsification. Recently, it has garnered significant research attention, with several studies investigating methodologies for converting pre-trained models into MoE architectures.
While various MoE approaches exist with different objectives, methods like Llama-MoE v2 focus on post-training optimization of instruction-tuned models, and approaches like DTSI target parameter efficiency during training from scratch. However, these methods either address specialized post-training scenarios or require training models from initialization, whereas our approach specifically targets sparsification of pre-trained dense models during the pre-training stage.
MoEficationZhang et al. (2022) trains routers to predict the activation patterns of experts that are partitioned from FFNs while keeping model parameters frozen, thereby activating a fixed number of experts. However, this method was primarily designed for ReLU activation functions and requires additional transformation steps for SiLU/GeLU activation functions that are widely utilized in contemporary Transformer architectures. FactorLLMZhao et al. (2024) employs a multi-stage training strategy, initially utilizing the original dense model to guide router training, followed by fixing the router and subsequently training the experts. This sequential training methodology constrains collaborative optimization between routers and experts, and its dependence on a teacher-student framework introduces additional training complexity. LLaMA-MoEZhu et al. (2024) explores the decomposition of FFNs and organizes training according to the Switch TransformerFedus et al. (2022) paradigm; however, it merely provides improved expert initialization while lacking flexible input-adaptive computation mechanisms. Given that MoEfication and FactorLLM differ significantly from mainstream MoE methods in architecture design and training paradigms, we choose to use LLaMA-MoE as a comparative approach.
Recent dynamic pruning methods such as DejaVuLiu et al. (2023) and PowerInferSong et al. (2024) can adaptively select activated weights based on input patterns. However, these approaches primarily focus on system-level acceleration through specialized hardware configurations: DejaVu requires integration with asynchronous hardware-aware implementations including kernel fusion and memory coalescing, while PowerInfer employs GPU-CPU hybrid inference engines to exploit locality patterns and minimize communication overhead. In contrast, our method employs algorithm-level sparsification, which reduces the model’s floating-point operations.
@@ 4 Method
Although our method is termed DSMoE, its training approach differs from traditional MoE methods such as Switch Transformer and DeepSeeKMoE Dai et al. (2024). Our objective is to achieve sparsity through partitioning pre-trained models, where each expert inherits a distinct portion of the original model’s knowledge. Our approach is based on the principle that the model should learn to selectively utilize different aspects of pre-trained knowledge based on input complexity, rather than routing tokens among independently trained experts. To implement this insight, we present our method in three modules.
@@ 4.1 FFN Partitioning
The widespread adoption of MoE architectures inspires our exploration of sparsity in FFN layers, suggesting that different parts of computation can be dynamically activated based on input patterns. Previous work has further revealed that FFN layers essentially operate as key-value memories, where different portions of the layer specialize in detecting and processing distinct input patternsGeva et al. (2020). Building on these insights, we propose to directly partition pre-trained FFN layers. As shown in Equation 5, we partition the matrices $\mathbf{U}$, $\mathbf{V}$, and $\mathbf{W}$ into $n$ groups along the intermediate dimension, where each group can be viewed as an “expert" that inherits a portion of the original transformation capabilities. When summing all expert outputs, this partitioned form is mathematically equivalent to the original FFN computation:
@@
To enable dynamic expert activation based on input, we employ a gating network that determines which experts should be activated. The expert’s output is propagated to the subsequent layer only when the corresponding gating activation value exceeds a certain threshold $\tau$. This can be formulated as:
@@
Model Configuration Params Activated Params PPL ($\downarrow$)
LLaMA-1B d=2048, D=8192 1.24B 1.24B 5.67
LLaMA-7B d=4096, D=11008 6.74B 6.74B 3.40
@@
LLM-Pruner-channel d=1215, D=8192 889M 889M 7.51
LLM-Pruner-block d=2048, D=3896.4 735M 735M 7.46
SparseGPT d=2048, D=8192 1.24B 735M 9.82
LLaMA-MoE d=2048, D=1024    $\times 8$, topK=3 1.24B 736M 7.45
DSMoE(ours) d=2048, D=1024    $\times 8$ 1.24B 735M 7.41
@@
LLM-Pruner-channel d=2401,   D=11008 3.95B 3.95B 4.01
LLM-Pruner-block d=4096,   D=6256.5 3.94B 3.94B 4.01
SparseGPT d=4096,    D=11008 6.74B 3.93B 3.96
LLaMA-MoE d=4096,    D=1376    $\times 8$, topK=3 6.74B 3.98B 4.12
DSMoE(ours) d=4096,   D=1376    $\times 8$ 6.74B 3.93B 3.91
Table 1: Results of perplexity (PPL) across different language models. The bold values indicate the best-performing method among various acceleration approaches. The Configuration column describes the specific model architecture, where $d$ represents the hidden dimension, D denotes the expansion dimension in FFN layers (for LLM-Pruner-block method, this represents the average value), $\times$ n indicates the use of n parallel FFN layers, and topK specifies the number of activated experts per layer in the MoE architecture. The Params column shows the total number of model parameters, while Activated Params indicates the average number of parameters activated during inference.
@@ 4.2 Straight-Through Estimator
This gradient blocking prevents non-activated experts from receiving training signals, leading to a “dead expert" problem where these experts become permanently inactive. Due to random initialization of the sigmoid gating parameters, experts with initially low activation probabilities below the threshold receive zero gradients and cannot improve through training, creating a Matthew effect where inactive experts remain progressively underutilized. Unlike traditional MoE models that train experts from scratch, our experts inherit pre-trained knowledge that we wish to preserve and adapt. To address this issue, we employ the straight-through estimator technique, which allows gradient flow through non-activated experts while maintaining thresholded activation during the forward pass:
S(x)=sg(G(x))+xsg(x)S(x)=sg(G(x))+x-sg(x) (10)
@@
where the operator $sg(\cdot)$" is the “stop gradient" operator to prevent gradient back propagation. The partial derivatives for experts and their gates below the threshold are as follows. Let:
$a_{i}=\text{act}(\hat{h}_{t}^{l}\mathbf{W}_{i}),\quad a^{\prime}_{i}=\text{act}^{\prime}(\hat{h}_{t}^{l}\mathbf{W}_{i})$ (12)
$g_{i}=\sigma(\hat{h}_{t}^{l}\mathbf{Y}_{i}),\quad u_{i}=\hat{h}_{t}^{l}\mathbf{U}_{i}$
@@
$\frac{\partial h_{t}^{l}}{\partial\mathbf{Y}_{i}}=(\hat{h}_{t}^{l})^{\top}\cdot(o_{i}\cdot\sigma^{\prime}(\hat{h}_{t}^{l}\mathbf{Y}_{i}))$
The gradient dynamics show a key property: with the straight-through estimator, experts receive gradients for their gating parameters regardless of activation status. The gradient direction for $\mathbf{Y}_{i}$ depends on whether the expert’s output $o_{i}$ would reduce the overall loss. This allows experts to adaptively learn when to activate based on their usefulness for specific input patterns.
@@ 4.3 Sparse Loss
$\mathcal{L}=\mathcal{L}_{\text{LM}}+\mathcal{L}_{\text{sparse}}$ (17)
where $\mathcal{L}_{\text{sparse}}$ denotes the sparsity loss term, which we abbreviate as $\mathcal{L}{\text{s}}$ in subsequent equations.
$\mathcal{L}=\mathcal{L}_{\text{LM}}+\frac{1}{LN}\sum_{l=1}^{L}\sum_{n=1}^{N}\mathcal{L}_{s}(G(\sigma(\hat{h}^{l}_{t}\mathbf{Y}_{n})))$ (18)
@@
$\mathcal{L}=\mathcal{L}_{LM}+\frac{1}{LN}\sum_{l=1}^{L}\sum_{n=1}^{N}G(\sigma(\hat{h}^{l}_{t}\mathbf{Y}_{n}))$ (19)
@@ 5 Experiments
Model Hellaswag LAMBADA PIQA SIQA StoryCloze Wino GSM8K NaturalQs TriviaQA WebQs
@@ 5.2 Experimental Setup
We evaluate DSMoE on two pre-trained models of different scales: Llama-7B [^1] https://huggingface.co/meta-llama/Llama-2-7b and Llama-1B [^2] https://huggingface.co/meta-llama/Llama-3.2-1B. For our method’s hyperparameters, we simply set the activation threshold $\tau=0.5$, learning rate to 2e-5, batch size to 32, and sequence length to 1024. To ensure fair evaluation, all baseline methods underwent continued training on identical data quantities (10B tokens) with the same training configurations.
We compare our approach with several baselines: the channel-wise and block-wise methods from LLM-Pruner (a structured pruning approach), and SparseGPT (an unstructured pruning method). To ensure fair comparison, all baseline methods (LLM-Pruner channel/block-wise, SparseGPT, and LLaMA-MoE) were trained on identical data quantities (10B tokens) and configured to match DSMoE’s activated parameter count.
Additionally, we compare against LLaMA-MoE, which applies a similar FFN partitioning scheme but follows the traditional MoE paradigm with fixed top-k expert selection and standard MoE training objectives, to investigate whether conventional MoE frameworks can effectively leverage pre-trained weights through warm-starting.
@@ 5.3 Main Results
We first present the model’s perplexity on the validation set. Following previous workTouvron et al. (2023); Brown et al. (2020); Su et al. (2024); Xiong et al. (2024); Dai et al. (2024), we then evaluate the model’s performance on downstream benchmarks, which includes zero-shot accuracy testing on HellaSwagZellers et al. (2019), LAMBADAPaperno et al. (2016), SIQASap et al. (2019), PIQABisk et al. (2020), StoryClozeMostafazadeh et al. (2016), and WinograndeSakaguchi et al. (2021). Additionally, we conduct 5-shot evaluation measuring exact match performance on TriviaQAJoshi et al. (2017), WebQuestions (WebQs)Berant et al. (2013), GSM8KCobbe et al. (2021), and Natural Questions (NaturalQs)Kwiatkowski et al. (2019).
@@ 5.3.1 Perplexity Results
Table 1 presents the perplexity results of the baseline dense model and its pruned, sparsified variants. The results demonstrate that DSMoE consistently outperforms baseline models under equivalent activation constraints. Our experimental results indicate that DSMoE achieves superior efficiency compared to static parameter pruning. Furthermore, DSMoE exhibits better performance than fixed-activation methods like MoE, which can be attributed to the fact that knowledge from all experts contributes to the model’s learning process, enabling it to develop the ability to flexibly select activations based on input. Additionally, DSMoE exhibits distinctive feature processing capabilities, learning layer-specific activation patterns that naturally emerge from the input complexity. We will examine these emergent patterns in detail in the analysis section.